OpenStudy (anonymous):

WHATS 6.25 AS A FRACTION

OpenStudy (cwrw238):

0.25 = 25/100 = 1/4 6.25 = 6 1/4 or 25/4

OpenStudy (anonymous):

6/1/4

OpenStudy (cwrw238):

its not usually written that way 6 1/4 is usual